OpenStudy (anonymous):

is f(x) = x^2 - 16/x-4 continuous at x = 4? Give reason for the answer.

OpenStudy (eyust707):

can you plug 4 in there and get a real number?

OpenStudy (anonymous):

^ ill plug a 4 into you....

OpenStudy (eyust707):

oh you know how much i love them 4s

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

\[(x ^{2} - 16)/(x-4) = [(x-4)(x+4)]/(x-4) = x+4\] So yes it is contiueous at x = 4 F(4) = 8

OpenStudy (anonymous):

First try plugging in 4 before you do any factorization. The function will be undefined ( i.e. division by zero) Next try to factorize (rearrange)the function. In this case (x^2 -16) = (x+4)(x-4) divide through by (x-4) and you are left with (x+4) The function is now defined at x =4. Now plug in 4 and it should be fine. lol

OpenStudy (anonymous):

Thank you very much

OpenStudy (anonymous):

vw

OpenStudy (anonymous):